OpenStudy (anonymous):

Find the slope of the straight line that passes through the given points or state that the slope is undefined. Then indicate if the line through the points rises (left to right), falls (left to right), is horizontal, or is vertical (1, 7) and (-2, -2)

OpenStudy (anonymous):

@iGreen @robtobey

OpenStudy (igreen):

Plug them into the slope formula. \(m = \dfrac{-2-7}{-2-1}\)

OpenStudy (anonymous):

the slope is always able to be found if you do what @iGreen is showing now! :D

OpenStudy (igreen):

Yep, and if the slope is positive, it'll rise from left to right. If it's negative, it'll fall from left to right.

OpenStudy (anonymous):

I answered this same question for him like 6 minutes ago xD

OpenStudy (anonymous):

Okay so it would be -2 - 7 = 5 and -2 - 1 = -1 ?

OpenStudy (igreen):

-2 - 7 isn't 5

OpenStudy (igreen):

-2 - 1 isn't -1 either

OpenStudy (anonymous):

Oh crap sorry. -2 -7 = 9 and -2 - 1 = -3 right?

OpenStudy (igreen):

-9, not 9

OpenStudy (igreen):

That gives us: -9/-3 Divide

OpenStudy (anonymous):

3 ?

OpenStudy (igreen):

Yep, that's the slope

OpenStudy (igreen):

So the line rises from left to right.

OpenStudy (anonymous):

Thank you for walking me through this. Your very helpful . Can I ask one more in another question box?

OpenStudy (igreen):

Sure thing